During 2018, Raines Umbrella Corp. had sales of $739,000. Cost of goods sold, administrative and selling expenses, and depreciation expenses were $454,000, $99,000, and $144,500, respectively. In addition, the company had an interest expense of $72,200 and a tax rate of 24 percent. (Ignore any tax loss carryforward provisions and assume interest expense is fully tax deductible.) a. What is the companys net income/loss for 2018? (Do not round intermediate calculations. Enter your answer as a positive value.) b. What is the companys operating cash flow? (Do not round intermediate calculations.) Answer is complete but not entirely correct. S 23,332 a. Net loss b. Operating cash flow
Answer #1

a) Net income = (Sales - COGS - Selling expenses - Depreciation - Interest expenses)*(1-Tax) =(739,000 -454,000-99000-144500-72200)*(1-24%) = -23332
Tax = (Sales - COGS - Selling expenses - Depreciation - Interest expenses)*Tax Rate = (739,000 -454,000-99000-144500-72200)*24% = -7368
Since EBIT is negative hence Tax = 0

b) OCF = EBIT -Taxes + Depreciation = (Sales - COGS - Selling expenses - Depreciation) - Taxes + depreciation =
(739,000 -454,000-99000-144500) -0 + 144500 = 186,000

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well